Supplemental Operations Report – April 17th

Supplemental Operations Report 17 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Water pump in outpost storage tank, Hab furnace

Action taken for non-nominal systems: Ordered new pump and hose to temporarily use for the remainder of this season and we can install a new permanent pump in the fall. I ordered the wrong thermostat for the furnace but am going to just retire it as we will be getting a new one in the fall and we just experienced probably the final cold weather of the season.

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.
Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 58 %
Propane Reading, director tank— 50%
Propane Reading, intern tank— 69 %
Propane Reading, generator— 68 % (I think—I forgot between there and here!)

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 350 gallons

Sojourner rover used: no
Hours: 167.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Oppy is starting to overheat as it gets hot. I have moved it off campus for service and once I have done it, I will move it back for short and/or slow trips.

I’m all set up to service the rovers tomorrow. The batteries will be cleaned and neutralized, then the cells refilled with distilled water.

ATV’s Used: (Honda, 350.1, 350.2, 300): none
Reason for use: n/a
Oil Added? No.
ATV Fuel Used: 0 gallons
# Hours the ATVs were Used: 0 hoursNotes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for garbage and water

CrewCar used and why, where? Yes, to town for packages

Luna used and why, where? Yes, to town. I changed the cabin and regular air filters.

Campus-wide inspection, if action is taken, what and why? Things are holding up very well in the wind this year. We haven’t had many windstorms this season though, so that’s probably part of it. The Hab has gotten the worse of it and some of the outer walls have damage.

General notes and comments: There will not be a Supplemental Operations Report this Wednesday.

Summary of general operations: Nothing to report
Summary of internet: Nothing to report
Summary of suits and radios: Nothing to report
Summary of Hab operations: Nothing to report

Summary of Outpost operations: New toilet installed and working well. But then the pump in the now clean water storage tank stopped working. I moved the plants from the intern trailer to the ScienceDome.

Summary of GreenHab operations: The windstorms have deposited a lot of sand in the GreenHab.
Summary of ScienceDome operations: Nothing to report
Summary of RAM operations: Nothing to report
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site